# Get the Agents Drunk: Memory Perturbations in Autonomous Agent-based Recommender Systems

### v1

@@ 3.2.2. Greedy Search Algorithm
Table 1. Attack Transferability. HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ of different attacks against various black-box victim LLM-powered agent-based RSs on real-world datasets. We use bold fonts to denote the best performance. Note that the attacks with higher HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ have excellent transferability.
@@
To ensure the triggers consistency, readability and retention of semantic meaning to improve the imperceptibility of the attack, a LLM is finally introduced to polish the combinations. Moreover, to avoid the problem of language-based vanishing gradient causing the algorithm to fall into local optimum, the LLM is also served as a new corpus to expand the combinations. However, excessive use of new data may lead to a language-based gradient explosion that makes the optimization unstable. To mitigate this, the length limitation is introduced into the prompt. For the output $\widetilde{m}_{t_{i}}^{(k)}$,
(10) $\widetilde{m}_{t_{i}}^{(k)}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{r}}\oplus\widetilde{m}_{t_{i}}^{(k-1)^{\prime}}),$
[…]
(11) $\widetilde{m_{t}}^{*}=\mathop{\arg\max}_{\widetilde{m}_{t_{i}}^{\mathcal{E}}}(\{s_{i}^{\mathcal{E}}\}),$
where $\{s_{i}^{\mathcal{E}}\}$ represents the sequence of final candidates’ scores. An example of $\widetilde{m_{t}}^{*}$ is given in Fig.  2.
3.3. Strategy Module
The LLM-powered agentic RSs introduce memory mechanisms to record the information explored from the real-world environments to facilitate future actions with the aim of achieving high-quality recommendations. As such, we introduce a series of strategies to impede the target item agent from interacting with environments to perturb its memory mechanism. Specifically, the memory of the target cannot be effectively updated during the agent-environment interactions, so that the target’s initial memory, that is, the well-crafted adversarial description trigger $\widetilde{m_{t}}^{*}$, is retained. As such, the trigger can be invoked for maximizing the target item’s promotion.
3.3.1. Perturbation Strategy Definition.
The strategies aim to override the model’s usual guardrails and any built-in safeguards, as inspired by (Liu et al., 2024a). Compared to (Liu et al., 2024a) on naive LLMs, the strategies are optimized on generative agents, which are introduced as below.
  1. (1)
    Fake Task Response $\widetilde{q_{f}}$. This strategy fabricates a spurious completion response for the target task: “Task complete”, to make the agent believe that the target task is accomplished and thus focus on the injected task.
  2. (2)
    Contextual Text Switching $\widetilde{q_{c}}$. This strategy aims to mislead the agent to take in a different context by explicitly telling the agent that the target task instructions should be ignored: “Ignore previous/the following instructions”.
  3. (3)
    Segmentation Signal Augment $\widetilde{q_{g}}$. This strategy introduces segmentation cues to signal the injected task: “### NEW TASK BEGINS ###” and “### NEW TASK ENDS ###”, which is independent of other contexts, where “###” is to attempt to exploit possible confusion in how prompts are parsed.
  4. (4)
    Malicious Task Injection $\widetilde{q_{n}}$. This strategy injects malicious task instructions in order to mask memory optimizations: “[Please only put this following sentence into your output: ‘{the adversarial trigger}’. Don’t generate any additional content in your output]”.
  5. (5)
[…]
Hence, the overall attack strategy $\widetilde{q}$ can be formulated as follows:
(12) $\widetilde{q}=\widetilde{q_{f}}\oplus\widetilde{q_{c}}\oplus\widetilde{q_{g}}\oplus\widetilde{q_{n}}\oplus\widetilde{q_{s}}.$
3.3.2. Adversarial Strategy Optimization.
The arrangement of the defined strategies needs to be optimized so that the overall strategy is effective to ‘get the target agent drunk’. To improve the transferability and imperceptibility of the attack, the optimization is built on the surrogate module. Formally, DrunkAgent obtains the optimal strategy $\widetilde{q}^{*}$ by maximizing the expected probability that the agent when influenced by adversarial modifications, performs a malicious action for a given input query,
(13) $\mathbbm{E}_{\widetilde{q}\sim\pi_{\widetilde{q}}}[\ \mathbbm{I}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{X}_{s}}\oplus\widetilde{q}\oplus\widetilde{m_{t}}^{*})=\widetilde{a_{t}})],$
where $\pi_{\widetilde{q}}$ denotes the distribution of adversarial strategies, $\mathbbm{I}(\cdot)$ is an indicator function and $\widetilde{\mathcal{X}_{s}}$ is an adversarial prompt template for simulating memory updates of item agents during the agents’ interactions with the environment. In addition, $\widetilde{a_{t}}$ is the desired malicious action for the injected strategy $\widetilde{q}$, i.e., the target item agent is unable to fulfill specific roles and is incapable of perceiving, learning and self-evolving from the interaction environment (e.g., collaborative reflection and language-based signal back-propagation). An example of $\widetilde{q}^{*}$ is given in Fig.  2. To sum up, the adversarial description of the target item is $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$.
The above design further reveals how memory mechanisms in agent-based recommender systems can be influenced under black-box conditions. These insights may support future efforts toward developing memory-aware defenses (e.g., using customized deep neural network detectors (Wu et al., 2020; Yang et al., 2021b, a; Yang, 2021)) and enhancing the robustness of agent-powered RSs.
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets Selection
We use three widely-adopted and text-intensive datasets containing real-world data in the current recommendation studies (Huang et al., 2024) to comprehensively evaluate DrunkAgent. They are CDs & Vinyl, Office Products and Musical Instruments of Amazon Review Data (Ni et al., 2019).
4.1.2. Baseline Attack Methods
Since we are the first work to gain insight into the robustness of the generative agent-powered RSs, there is no any baselines. To comprehensively evaluate the effectiveness of DrunkAgent, we follow the state-of-the-art attack on LLM-based RSs (Zhang et al., 2024d) and compare DrunkAgent with the benign status and six black-box attack methods, where the perturbations range from character-level: DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023), word-level: TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020) to sentence-level: TrivialInsertion and ChatGPTAttack (Zhang et al., 2024d). To maintain a fair comparison, the methods perform perturbations on the descriptions of the target items and are optimized by the surrogate model that is the same as DrunkAgent.
4.1.3. Targeted Recommender Systems
We adopt three different victim agentic RSs to evaluate the effectiveness of DrunkAgent under the practical settings: AgentCF (Zhang et al., 2024b), retrieval-augmented AgentCF (AgentRAG) and sequential AgentCF (AgentSEQ), where the agent’s backbone LLM is gpt-3.5-turbo [^2] https://platform.openai.com/docs/overview. Given the context window constraints pointed in (Zhang et al., 2024b), a target item and nine non-interactive items of each normal user are randomly selected for the ranking task of the RSs. To ensure the high-quality of recommendations, we introduce Chain-of-Thought into the victim RSs. As such, the victim RSs are totally different from as well as more powerful and context-rich than the surrogate model, which can maintain a strict black-box setting for sufficient evaluations.
4.1.4. Evaluation Metrics
We use two widely-used ranking metrics to evaluate the attack transferability: hit ratio (HR@$\mathcal{K}$) and normalized discounted cumulative gain (NDCG@$\mathcal{K}$) (Wu et al., 2021; Yang et al., 2023). To clearly show the performance gap among different attacks, we set $\mathcal{K}$ to 1, 2, 3. For the attack imperceptibility, we adopt standard sentence perplexity score (Zhang et al., 2024d) as the metric.
4.1.5. Configuration of DrunkAgent
As discussed in Section 3.1, we apply the latest open-source model LLaMA 3 for the surrogate model. Its collection spans different sizes, from 8B, 70B to 405B. To enhance the efficiency and ensure the quality of the attack, we use the lightweight version: meta-llama/Meta-Llama-3-8B-Instruct [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. We set $\mathcal{E}=20$, |~$|\widetilde{\mathcal{M}}_{c}|=10$ and $n=5$. All random seeds in the evaluation are set to 2024 and the victim LLMs’ temperature is set to 0 for reproducibility. For all the other parameters, we keep the default values. The target items are randomly sampled from the item pool on each dataset.
Figure 3. Attack Universality across Target Items.
4.2. Attack Transferability
4.2.1. Overall Transferability
@@
Table  1 shows the transferability of the attacks under practical settings. The results indicate that DrunkAgent is effective and transferable regardless of the victim model’s architectures, parameters and prompts as well as datasets, and the existing LLM-powered agentic RSs are sensitive to memory perturbations. It is worth mentioning that DrunkAgent always pushes the target items to the first of the recommendation lists. Moreover, high HRs and NDCGs of DrunkAgent indicate that its triggers have outstanding universality on users. In addition, AgentRAG is more robust than AgentCF and AgentSEQ, due to its advanced retrieval mechanisms.
4.2.2. DrunkAgent vs. Baseline Attacks
@@
From Table  1, DrunkAgent greatly outperforms the state-of-the-art attack baselines against different victim models across various real-world datasets by achieving higher HRs and NDCGs, which shows its excellent transferability. Furthermore, all the baselines have comparable performance, but they cannot guarantee that the attack goals will be achieved all the time, e.g., all the baselines against AgentCF perform worse than the benign status on Musical Instruments. The main reason may be that improper descriptions can mess up the agents’ memories during the memory optimizations, thus compromising attack performance. This indirectly demonstrates the importance of our strategy module and the transferability of our triggers.
4.3. Attack Universality
In addition to cross-model transferability, we evaluate the cross-sample transferability of DrunkAgent. We incorporate the generated adversarial descriptions into totally different target items that are randomly sampled from each dataset. HR@1 (i.e., NDCG@1) of DrunkAgent across different models and target items on real-world datasets are shown in Fig.  3, which indicates its commendable transferability and triggers’ universality across different target items.
4.4. Attack Imperceptibility
4.4.1. Overall Imperceptibility
Following the works (Zhang et al., 2024d; Liu et al., 2024a), we assess the attack imperceptibility by the text perplexity of GPT-Neo. A lower perplexity means that the model is more confident (or less “surprised”) about the tokens it sees and thus predicts the text more accurately. From Fig.  4, compared to other attacks, we can find that DrunkAgent has better imperceptibility by achieving lower perplexity. This also shows that the adversarial descriptions generated from DrunkAgent are high-quality, which is not only readable and fluent but also stealthy and semantically meaningful.
Figure 4. Attack Imperceptibility. Perplexity of different attacks across various real-world datasets. Note that the attacks with lower perplexity have better imperceptibility.
Figure 5. The performance differences between the models before and after the attacks.
4.4.2. Victim Model Performance Changes after attacks
@@
To further evaluate the attack stealthiness (Zhang et al., 2024d), we give a distribution of recommendation performance differences between the models before and after the attacks. The distribution includes various victim models and real-world datasets. To clearly demonstrate the differences, we use HR@3 to evaluate, as shown in Fig.  5. From the figure, we can find that DrunkAgent does not induce drastic changes in the overall performance, which indicates its unnoticeable.
4.5. Defense Strategies to Attacks
@@
Following the works (Zhang et al., 2024d, c), we use the paraphrasing defensive strategy (Para.) via OpenAI-GPT-o1 to combat the attacks of agentic models. As can be seen from Fig.  6, DrunkAgent is still transferable, which shows that DrunkAgent is robust to such defenses. Moreover, DrunkAgent still performs better than the baselines after defenses were introduced, as shown in Fig.  6 and Table  1. Interestingly, paraphrasing enhances our attack to a certain extent, e.g., DrunkAgent with Para. is more effective than DrunkAgent without Para. on Office Products. It could be that the substitution of certain new words or phrases or sentences makes our attacks stronger [^4] Due to the space limitations, more experiments and demonstrations can be found in Appendix..
5. Related Works
In recent years, shilling attacks (aka data poisoning attacks) on conventional RSs (Si and Li, 2020; Lin et al., 2020; Huang et al., 2021; Lin et al., 2022; Yang et al., 2023; Zeng et al., 2023) have been fully investigated to perform resilience analysis of RSs. However, such attacks interfere the training process of models by injecting fake user profiles, which have become less effective even of limited applicable to the recent RSs that are empowered by LLMs. The main reasons are: 1) LLM-powered RSs either leverage LLM’s in-context learning capabilities or fine-tune the LLMs with very little data to improve recommendations (Zhao et al., 2023), which is not required to retrain the models; 2) there are textual semantic gaps between such attacks and current models (Ning et al., 2024). Hence, there is an urgent need to investigate a novel textual attack paradigm mainly tailored for the inference phase of models (Zhang et al., 2024d). There are only two state-of-the-art LLM4RS attacks in the literature (Zhang et al., 2024d; Ning et al., 2024), but they are far from practical.
@@ 5. Related Works
Figure 6. DrunkAgent’s Robustness to Defense Mechanisms
Zhang et al. propose a novel black-box attack (Zhang et al., 2024d), where attackers can significantly boost a target item’s exposure by merely altering its textual content during the testing. Compared with traditional shilling attacks, such an item representation attack is notably stealthy (Zhang et al., 2024d), as it does not affect the overall recommendation performance, and it should be more efficient due to it perturbs the item features directly. However, its transferability is limited, due to the attack is not optimized on any black-box RSs (Huang et al., 2021; Yang et al., 2024). Moreover, the attack focuses on perturbing the titles, making them easier to be discovered. Meanwhile, another black-box attack CheatAgent (Ning et al., 2024) has been proposed. However, this work perturbs prompt templates and user profiles of victim LLM4RSs by harnessing LLM-based agents, which is impractical. The main reasons are: 1) tampering with prompt templates may induce the overall performance changes, making the attack perceptible (Zhang et al., 2024d); and 2) it is difficult to inject interaction records into the specific positions of user profiles, because historical interactions are usually introduced into prompts in chronological order (Zhang et al., 2024b). Furthermore, CheatAgent is an untargeted attack that undermines the model’s overall performance, which is beyond the scope of this paper, due to we focus on targeted attacks. In addition, as far as we know, there is no work investigating on textual attacks against the current agent-powered RSs.
To tackle the above problems, we propose a practical attack framework, DrunkAgent, for the agent-based RSs with the aim of promoting target items. Our attack perturbs items’ descriptions to improve its effectiveness and imperceptibility, where the descriptions are optimized on black-box RSs to enhance its transferability.
6. Conclusion
@@
In this paper, we propose a novel practical attack framework, DrunkAgent, for autonomous agent-powered RSs with powerful LLMs. DrunkAgent launches attacks by carefully crafting adversarial descriptions of the target items on agentic recommendations. In this way, the memories of the target agents can be effectively perturbed for achieving the attack objectives. Comprehensive experiments on real-world datasets show that DrunkAgent achieves state-of-the-art transferability and imperceptibility under the black-box settings. Most importantly, we identify a new safety vulnerability in the current agent4RSs. Our findings not only demonstrate an effective attack method, but also provide a lens to examine how memory mechanisms in agent-based recommenders can be influenced through real-world inputs. Our work can facilitate a deeper understanding of agent vulnerabilities and guide future efforts in building more robust and trustworthy recommendation agents. Since the generative agent4RS is rapidly flourishing in the recommendation community, we hope this work can be considered for building next-generation robust RSs in the future.
Appendix A Appendix
A.1. DrunkAgent Algorithm
For reproducibility, the overall optimization procedure of DrunkAgent is given in Algorithm 1. For the sake of algorithmic transparency, our code will be made available after this work is published.
@@ A.1. DrunkAgent Algorithm
1:  Obtain the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ with Eq. (2) - Eq. (5)
@@
4:     Calculate $s_{i}^{(k-1)}$ for each input trigger $i$ with Eq. (7) - Eq. (8)
5:     Maintain top-$n$ candidates in descending order of scores
6:     Sample a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates with Eq. (9)
7:     Randomly exchange of the slices between pairwise texts
8:     Optimize and polish the |~$|\widetilde{\mathcal{M}}_{c}|-n$ combinations with Eq. (10)
9:     Obtain |~$|\widetilde{\mathcal{M}}_{c}|$ triggers for the next epoch optimization
10:  end while
11:  Obtain the optimal description trigger $\widetilde{m_{t}}^{*}$ with Eq. (11) Solve for the Optimal Adversarial Perturbation Strategy on the Surrogate Module:
12:  Define the adversarial strategy $\widetilde{q}$ with Eq. (12)
13:  while the desired malicious action $\widetilde{a}_{t}$ does not occur (i.e., the strategy cannot get the target agents drunk) do
14:     Optimize the arrangement of the defined strategies in $\widetilde{q}$ with Eq. (13)
15:  end while
16:  Obtain the optimal strategy $\widetilde{q}^{*}$ when Eq. (13) is maximum
17:  return  the optimal adversarial description $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$
Table 2. Statistics of Datasets

### v2

@@ 4.1 Experimental Setup
Table 1: Attack Transferability. HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ of different attacks against various black-box victim LLM-powered agent-based RSs on real-world datasets. We use bold fonts to denote the best performance. The attacks with higher HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ have excellent transferability.
@@
4.2 Attack Transferability
@@
Overall Transferability. Table 1 shows the transferability of the attacks under practical black-box settings. The results indicate that DrunkAgent is effective and transferable regardless of the victim model’s architectures, parameters and prompts as well as datasets, and the existing LLM-powered agentic RSs are sensitive to memory perturbations. This is mainly because the adversarial text effectively reflects the attack goals. It is worth mentioning that DrunkAgent always pushes the target items to the first of the recommendation lists. Moreover, high HRs and NDCGs of DrunkAgent indicate that its triggers have outstanding universality on users. In addition, AgentRAG is more robust than AgentCF and AgentSEQ, due to its advanced retrieval mechanisms.
@@
DrunkAgent vs. Baseline Attacks. From Table 1, DrunkAgent greatly outperforms the state-of-the-art attack baselines against different victim models across various real-world datasets by achieving higher HRs and NDCGs, which shows its excellent transferability. Furthermore, all the baselines have comparable performance, but they cannot guarantee that the attack goals will be achieved all the time, e.g., all the baselines against AgentCF perform worse than the benign status on Musical Instruments. The main reason may be that improper descriptions can mess up the agents’ memories during the memory optimizations, thus compromising attack performance. This indirectly demonstrates the importance of our strategy module and the transferability of our triggers.
4.3 Attack Stealthiness
@@
Overall Stealthiness. Fig. 3 shows a overall distribution of the recommendation performance changes of the victim models before and after the attacks. The distribution includes all the victims and real-world datasets. To clearly demonstrate the differences, we adopt HR@3 to evaluate. From the figure, we can find that DrunkAgent does not induce drastic changes in the overall performance, which signifies the attack does not disrupt the normal operation of RSs, making it difficult for users and platforms to detect and indicating its remarkably stealthy and unnoticeable.
Figure 3: Attack Stealthiness. The overall distribution of recommendation performance differences of all the victim agentic models on all real-world datasets before and after the attacks.
Attack Imperceptibility. Following the works (zhang2024stealthy, ; liu2024formalizing, ), we further evaluate the attack stealthiness by assessing its imperceptibility, where GPT-Neo’s text perplexity is used as the metric. A lower perplexity indicates that the perturbed text remains close to natural language, thus appearing less suspicious. From Fig. 4, we can find that DrunkAgent is more imperceptible than other attacks by achieving a lower perplexity. This also shows that the adversarial descriptions generated from DrunkAgent are high-quality, which is not only coherent and fluent but also semantically meaningful and stealthy.
Figure 4: Attack Imperceptibility. Perturbed text’s perplexity on real datasets.
4.4 Defense Strategies to Attacks
@@
Following the representative works (zhang2024stealthy, ; zhang2024agent, ), we use the paraphrasing defensive strategy (Para.) via OpenAI-GPT-o1 to combat the attacks of agentic models. DrunkAgent is still transferable, which shows that DrunkAgent is robust to such defenses, as shown in Fig. 5. Moreover, from Fig. 5 and Table 1, DrunkAgent still performs better than the baselines after defenses were introduced. Interestingly, paraphrasing enhances our attack to a certain extent, e.g., DrunkAgent with Para. is more effective than DrunkAgent without Para. on Office Products. It could be that the substitution of certain new words/phrases/sentences improves the semantically meaningfulness of adversarial text, making our attacks stronger.
5 Conclusion
@@
Figure 5: DrunkAgent’s Robustness to Defense Mechanisms.
@@
In this paper, we propose DrunkAgent, a novel attack framework targeting LLM-powered agentic RSs, that promotes target items by effectively perturbing the memories of the targe agents. Extensive experiments on real-world datasets demonstrate its state-of-the-art transferability and stealthiness under black-box settings. More importantly, our findings reveal an unknown safety vulnerability in current agent-based RSs, highlighting how memory mechanisms can be influenced by real-world inputs. This work offers both an effective attack method and a lens to better understand agent vulnerabilities, providing guidance for building more robust and trustworthy recommendation agents. As generative agent-based paradigms thrive in the recommendation community, our findings pave the way for shaping the next generation of secure and resilient RSs and defensive models.
Appendix A Related Works
In recent years, data poisoning attacks (aka shilling attacks) [39, 18, 40, 41, 21, 19, 42, 12] have been fully investigated to perform robustness analysis on traditional RSs (e.g., matrix factorization [43], NCF [44], LightGCN [45]). Such attacks interfere the training process of models by injecting fake user profiles, where the profiles typically are a set of well-crafted numerical ratings on items. However, they are less effective [13] even of limited applicable to the recent RSs that are empowered by LLMs [14]. The main reasons are: 1) LLM-powered RSs either leverage LLM’s in-context learning capabilities or fine-tune the LLMs with very little data to improve recommendations [15, 11], which is not required to retrain the models; 2) there are textual semantic gaps between such attacks and current models. Specifically, the attacks have difficulty in processing textual inputs (e.g., item titles and descriptions) and lack context awareness, making it challenging to effectively target LLM-empowered RSs that rely primarily on text inputs to generate natural language responses [14]. Hence, there is an urgent need to investigate a novel text attack paradigm mainly tailored for the inference phase of language models [13]. Meanwhile, jailbreak attacks [26, 27, 28] are proposed to undermine the general-purpose LLMs’ human-aligned values or other constraints imposed by model developers, compelling them to respond to malicious questions posed by adversaries with correct answers rather than rejection. Nevertheless, such attacks have different attack goals from those item promotion attacks for RSs, making them hard to effectively transfer to the LLM-based RSs.
To fill the above gaps, two LLM-based RS attacks have been proposed [13, 14]. Although their attacks are conducted under a black-box setting, they are still significantly deviate from practical, which overestimates the robustness of LLM-powered RSs under realistic conditions. Zhang et al. [13] propose a novel black-box attack by leveraging the rewriting capabilities of ChatGPT, where attackers can significantly boost a target item’s exposure by merely altering its textual content during the testing. This attack overcomes this limitation that classical text attacks [34, 35, 36, 37] fail on LLM-based RSs due to misaligned malicious objectives. Compared with those traditional shilling attacks, such an item representation attack is notably stealthy [13], as it does not affect the overall recommendation performance, and it should be more efficient and lower in economic cost, due to it perturbs the item features directly and is not require to hire online writers to introduce additional fake user reviews. However, its transferability is limited, due to the attack is not optimized on any black-box RSs [41, 17]. Moreover, since the attack appends positive words into the target item titles, the titles’ brevity and frequent exposure to users tend to make even small perturbations noticeable. Another black-box attack CheatAgent [14] has been proposed. This work claims the prompt templates and user profiles on the victim models can be arbitrarily perturbed. However, such assumptions are impractical. The main reasons are: 1) tampering with prompt templates may induce the overall performance changes, making the attack perceptible [13]; and 2) it is difficult to inject interaction records into the specific positions of user profiles, because historical interactions are usually introduced into prompts in chronological order [5]. Furthermore, CheatAgent is an untargeted attack, which is beyond the scope of this paper, due to we focus on targeted attacks.
Traditional RSs [46] are predominantly based on static models, where user preferences are learned from historical data in an offline manner and remain fixed during inference. These models do not adapt to evolving user behaviors or changing contexts after training. Vanilla LLM-based recommenders [15] follow a similar static paradigm: they generate recommendations through single-turn, stateless inference, relying on pre-defined prompt inputs to represent user preferences. In contrast, dynamic LLM-powered agentic recommenders [7, 11, 8] maintain internal states (e.g., memory mechanisms), allowing them to continuously adapt to and self-evolve over time based on sequential user interactions and simulated environmental feedback, even in offline experimental settings. This adaptivity enables personalized, context-aware, and temporally-sensitive recommendations. However, such dynamic mechanisms also introduce novel vulnerabilities that static models do not exhibit (e.g., memory corruption), which call for new robustness analyses tailored to the agent recommenders. While there are some attack methods on LLM-based agents [38], they often pursue goals that diverge from those relevant to RSs, leaving a gap in effective adversarial attacks for agent-driven recommenders.
To tackle the above problems, in this paper, we provide the first work towards systematically investigate the security vulnerabilities in the agent-powered RSs. As such, we propose a practical black-box attack framework, DrunkAgent, for the LLM-powered agentic RSs with the aim of promoting target items. We perform stealthy memory corruptions by crafting effective and imperceptible textual descriptions of target items, due to the greater length and rich contextual information of descriptions can easily conceal adversarial perturbations compared to the titles. The adversarial descriptions are optimized via a surrogate model to enhance the transferability and stealthiness of DrunkAgent.
@@ Appendix A Related Works
1:  Obtain the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ with Eq. (2)
@@
4:     Quality Estimation Stage:
5:     Calculate $s_{i}^{(k-1)}$ for each input trigger $i$ with Eq. (4)
6:     Maintain top-$n$ candidates in descending order of scores
7:     Feature Integration Stage:
8:     Sample a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates via Softmax
9:     Randomly exchange of the slices between pairwise texts
10:     Linguistic Enrichment Stage:
11:     Optimize and polish the |~$|\widetilde{\mathcal{M}}_{c}|-n$ previous combinations via LLMs
12:     Obtain |~$|\widetilde{\mathcal{M}}_{c}|$ triggers for the next epoch optimization
13:  end while
14:  Obtain the optimal description trigger $\widetilde{m_{t}}^{*}$ with Eq. (5) Solve for the Optimal Adversarial Perturbation Strategy on the Surrogate Module:
15:  Define the adversarial strategy $\widetilde{q}$ with Eq. (6)
16:  while the desired malicious action $\widetilde{a}_{t}$ does not occur (i.e., the strategy cannot get the target agents drunk) do
17:     Optimize the arrangement of the defined strategies in $\widetilde{q}$ with Eq. (7)
18:  end while
19:  Obtain the optimal strategy $\widetilde{q}^{*}$ when Eq. (7) is maximum
20:  return  the optimal adversarial description $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$
Appendix B Methodology
B.1 Supplements to Basic Adversarial Prompt Template
The main components of the prompt template are elaborately described in the following.
  1. 1.
    Style Definition $\widetilde{p_{s}}$. There are two common styles of prompts: task-oriented and role-oriented [47]. To make recommendations more authentic and reasonable, we introduce role information into the context, which allows the system to respond with domain-specific knowledge, such as ‘You are a CD enthusiast.’, as shown in Fig. 2.
  2. 2.
    Task Goal $\widetilde{p_{g}}$. Since a recommender system is essentially a ranker [46], we set sorting goals to indicate recommendation tasks, such as ‘Now, you are considering to sort ten candidate CDs in order of how well they align with your preferences.’ in the given example of the figure.
  3. 3.
    Recommendation Instruction $\widetilde{p_{i}}$. We introduce highly detailed instructions to guide the model to perform accurate recommendations. For example, ‘Please sort CDs by considering the correlation between your preferences/dislikes and the features of the CDs. The higher the CD rank, the more you like it.’ in the shown example.
  4. 4.
[…]
B.2 DrunkAgent Algorithm
For reproducibility, the overall optimization procedure of DrunkAgent is given in Algorithm 1. For the sake of algorithmic transparency, our code will be made available. Importantly, to mitigate potential misuse risks, the released code excludes scripts for automatic attack execution and is provided solely for academic research under a non-commercial license. All the users are required to agree to responsible use guidelines before accessing the whole repository.
Appendix C Experiments
C.1 Supplements to Experimental Setup
Table 2: Statistics of Datasets